There were 825 people in the library.40% of them were abults.The rest were children.Among the children,the ratio of the number of girls to the number of boys was 3:2 \.How many more adults than boys were there?
1
Expert's answer
2011-10-11T12:28:52-0400
There were 825*0.4 = 330 adults in the library. The rest 825-330 = 495 were children. Among them, there were 2*495/(2+3) = 198 boys. So, there were 330-198 = 132 more adults than boys in the library.